OpenStudy (anonymous):

Misha increased her test score from 60% to 80%. What was the approximate percent increase of Misha's test score?

OpenStudy (anonymous):

choices

OpenStudy (anonymous):

25% 33% 66% 75%

OpenStudy (anonymous):

ok so what is 60 divided by 80

OpenStudy (anonymous):

0.75

OpenStudy (anonymous):

so wat is 75% - 100%

OpenStudy (anonymous):

i hope you didnt put that as the answer

OpenStudy (tkhunny):

I would guess the percents are confusing. Increase from 60 to 80. Increase is 80 - 60 = 20 20 is what percent of 60?

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

cause its not out of 80 its out of 100

OpenStudy (tkhunny):

I told you the percents were confusing. 20 is what % of 60?

OpenStudy (anonymous):

30 percent but it is a test wich the highest you can score is a hundred so its based of that

OpenStudy (tkhunny):

30% is incorrect. 100 means nothing. If I asked this, what would be the result? I used to have 60 sheep. Now I have 80 sheep. What was the percent increase in my sheep population?

OpenStudy (anonymous):

but are percents are out of a hundred so thats not the same case at all

OpenStudy (tkhunny):

It is exactly the same. This is why I said the percents might be confusing. 100 is still nothing. We had 60% We changed to 80% What was the increase? 80% - 60% = 20% What was the percent increase? 20% / 60% = 1/3 = 0.33333... = 33.3333% We had 60 sheep We changed to 80 sheep What was the increase? 80 sheep - 60 sheep = 20 sheep What was the percent increase? (20 sheep) / (60 sheep) = 1/3 = 0.33333... = 33.3333% EXACTLY the same.

OpenStudy (anonymous):

its 25% i have had this question and got it right

OpenStudy (tkhunny):

If that is so, and I have no reason to believe it is not as you have stated it, then the answer is incorrect. It is as I have stated it. 60 ==> 80 is a 33.333% increase. 80 ==> 60 is a 25% decrease. It is a common point of confusion. You should challenge the author of the test and get this incorrect answer repaired. I suspect many other students have been misled by it.